ideal transformer in Fig. 9 has 500 turns on the ary. The source produces a voltage Eg of 600 V, and the load Z is a resistance of 12 . Calculate the following: a. The voltage E2 b. The current /2 c. The current I d. The power delivered to the primary [W e. The power output from the secondary WT In Problem 4, what is the impedance seen by the source Eg? primary and 300 turns on the second- 5

media%2Fae8%2Fae85c28d-080f-4de6-aaca-9d

I only need problem 5 please help I think I know how to do it but just want to be sure.
Answer #1

5) Impedance seen by source Eg is

Z'=Z \times \left ( \frac{N_{_{1}}}{N_2} \right )^{2}

=12 \times\left ( \frac{500}{300} \right )^{2}

= 33.33 \Omega

Thus, the impedance seen by the source Eg is 33.33\Omega